Bob has $1000. He decides to put his money in a Kenyan bond dominated in Kenyan shillings (K) that matures in a year. If the dollar-shilling spot exchange-rate decreases the day before Bob converts his dollars into shillings (i.e., the dollar appreciates), will the dollar return on his investment be higher or lower compared to the estimated dollar return before the exchange-rate change? Explain how you reached your answer. Assume there is no cost to exchanging currency. (Hint: you need to use the asset approach to exchange-rate determination.)
Answer #1

The dollar return on his investment is lower compared to the estimated dollar return prior to the exchange-rate change
the above is the response. because the greater worth of the dollar will minimize your return of investment in other currencies at the time of conversion into dollar
I will discuss the above using the below example
assume current dollar-shilling spot exchange-rate is that $1= 100 shillings, so you invested in Kenyan bond, dominated= 1000 * 100= 100000.
now state you made 10% on Kenyan bond in return as interest income, so your worth of the investment in the shilling.
= 100000 *( 1 +10%)= 110000 and return is 10% in shilling.
now say that dollar has actually appreciated and new dollar-shilling spot exchange-rate is that $1= 120 shilling.
Now transform your shilling financial investment in dollars which will be= 110000/120= 916.67 in US dollars.
Return in dollar terms= 916.67/ 1000-1= -8.33%.
so you can see dollar returns are lower than shilling returns due to gratitude in the dollar value.
the above will be addressed.